# Quaternion Multi-focus Color Image Fusion

### v1

@@ IV-F Ablation study
Parameter $\bm{\alpha=0}$ 1.0580 0.7359 0.7488 0.8240 0.7298
$\bm{\beta=0}$ 1.0052 0.7338 0.7282 0.8279 0.6980
$\bm{\lambda=0}$ 1.0140 0.7324 0.7432 0.8259 0.7127
QSD 1.0685 0.7374 0.7832 0.8234 0.7508
Patch size $\bm{5\times 5}$ 1.0503 0.7376 0.7277 0.8263 0.7115
@@
Fusion QBDF-Base-scale 0.9312 0.7021 0.7113 0.8099 0.6830
@@
Table III gives the quantitative fusion results with different patch sizes of the QSD model. As can be seen, the size $8\times 8$ provides the best quantitative results among five metrics.
Effectiveness of QBDF and QSSR. To verify the advantages of the QBDF and the QSSR strategies, we perform the ablation experiments on the mffw data. This includes the first 10 pairs of partially focused color images under various exposures and obvious misalignment due to severe defocus spread effects. First, the coefficient matrix and the detail layer of quaternion representation of each input color image are extracted using Algorithm 1. Then, base-scale and detail-scale fused results are obtained using the focus measures from Algorithm 2. These results are subsequently combined to generate the final output using the QSSR strategy.
Figure 8: Visual comparison of the background-focused input and various fused results including QBDF-Base-scale, QBDF-Detail-scale and QSSR results. The first row represents the input and fused results. The second row represents the differences between the fused results and corresponding background-focused input.
Fig. 8 provides visual comparison of the background-focused input color image and three fused results. As can be observed, QBDF-Detail-scale fusion introduces noticeable artifacts in the wall regions and creates unrealistic boundary lines within the yellow bounding box. In contrast, the QBDF base-scale fusion achieves smoother integration in low-texture areas but fails to preserve fine details in structured regions such as piano keyboards, resulting in oversmoothed edges. QSSR addresses these problems and significantly improves fusion performance in low-gradient, small-area focused, and boundary regions.
The analysis of the QBDF and the QSSR strategies is also quantitatively presented in this ablation study. Table III gives the quantitative fusion results with different scales of measures. As can be observed, QSSR provides the best quantitative results among five metrics. To verify the robustness of our QSSR strategy, The weight $\tau$ of QSSR is discussed here. There are two different settings for $\tau$ in Eq. (17). The adaptive weight $\tau$ in Eq. (18) provides the best quantitative results among five metrics.
Effectiveness of QMCIF. Table III also evaluates the impact of different representation strategies on fusion performance by comparing our QMCIF framework with conventional real-domain fusion method. To separate the benefits of quaternion representation, we replicate the fusion process in the real domain by converting color images to grayscale. Quantitative results demonstrate that the quaternion representation achieves superior performance across all metrics, particularly outpeforming real domain-based framework by $12\%$ in $Q_{CB}$.
@@ V Conclusion
We proposed a quaternion multi-focus color image fusion framework to perform high-quality color image fusion completely in the quaternion domain. This framework is flexible to be extended from fusing two color images at a time into fusing multiple color images simultaneously under complex challenging scenarios. This framework proposed a quaternion sparse decomposition model for high-precision focus detection of color images. This model jointly learns a low-rank structured coefficient matrix and a detail-scale layer. They both were utilized to estimate the focus level of the image. To solve the optimization problem of the QSD model, we developed an iterative algorithm under the framework of the quaternion alternating direction methods of multipliers. A quaternion base-detail fusion strategy was also introduced for high-quality base-scale and detail-scale color image fusion individually. A quaternion structural similarity strategy was further introduced to balance the trade-off between focus information preservation and artifact removal. Our framework fuses two-scale quaternion matrices and compares them with the input quaternion representations to adaptively select the focused patches with a higher QWSSIM value to form the final fused results. Extensive experiment results have shown that our framework outperforms the state-of-the-art methods in terms of focus information preservation and artifact removal under real-world challenging scenarios.

### v2

@@ IV-D Ablation study
Patch size for QCAFD $\bm{5\times 5}$ 1.0503 0.7376 0.7277 0.8263 0.7115
@@
Fusion for QBDF and QSSR QBDF-base-scale 0.9312 0.7021 0.7113 0.8099 0.6830
@@
Impact of patch size in QCAFD. Table IV shows the quantitative analysis of different patch sizes within the QCAFD module. For QCAFD, we try different patch sizes to perform the QFED process and the patch-wise base-scale focus map generation for QBDF and QSSR strategies. The optimal performance is achieved with a patch size of $8\times 8$, demonstrating balanced sensitivity to both high-texture and low-gradient regions.
Effectiveness of QBDF and QSSR. To verify the advantages of the QBDF and QSSR strategies, we perform the ablation experiments on the mffw data. This includes the first 10 pairs of partially focused color images under various exposures and obvious misalignment due to severe defocus spread effects. First, the coefficient matrix and the detail layer of quaternion representation of each input color image are extracted using Algorithm 1. Then, base-scale and detail-scale fused results are obtained using the dual-scale focus measures. These results are subsequently combined to generate the final output using the QSSR strategy.
Figure 9: Visual comparison of the background-focused input and various fused results including base-scale fusion, detail-scale fusion and the QSSR strategy. The first row represents the input and fused results. The second row represents the differences between the fused results and corresponding background-focused input.
The importance of the QBDF and QSSR strategies is highlighted through qualitative and quantitative comparisons. Fig. 9 visually demonstrates while detail-scale fusion alone leads to artifacts and base-scale fusion smooths out fine details excessively. The QSSR strategy effectively preserves important structural details and eliminates artifacts, especially in boundary and low-gradient regions. The quantitative results in Table IV are align with this visual assessment. This indicates the improvements across multiple metrics when employing QSSR.
Adaptive weighting in QSSR. We further evaluate the robustness of adaptive weighting in the QSSR strategy (Eq. (21)). Table IV reveals that adaptive weighting consistently outperforms fixed weighting across all evaluated metrics. This confirms that the adaptive weighting accurately assesses local structural similarity to optimize patch selection.
Effectiveness of QMCIF. Table IV compares the quaternion representation with an equivalent real-domain under our framework. We replicate the fusion process in the real domain by converting color images to grayscale. Quantitative results in Table IV highlight inherent advantages of quaternion representation in preserving inter-channel correlations and structural coherence during fusion.
@@ V Conclusion
This paper proposed a quaternion multi-focus color image fusion framework to perform high-quality color image fusion completely in the quaternion domain. This framework is flexible to be extended from fusing two color images at a time into fusing multiple color images simultaneously under complex challenging scenarios. This framework proposed a quaternion consistency-aware focus detection method. It consists of a quaternion focal element decomposition (QFED) module that jointly learns a low-rank structured coefficient matrix and a detail-scale layer and a dual-scale focus map generation strategy for robust focus detection. To solve the optimization problem of the QFED model, we developed an iterative algorithm under the framework of the quaternion alternating direction methods of multipliers. A quaternion base-detail fusion strategy was also introduced for high-quality base-scale and detail-scale color image fusion individually. A quaternion structural similarity strategy was further introduced to balance the trade-off between focus information preservation and artifact removal. Our framework fuses two-scale quaternion matrices and compares them with the input quaternion representations to adaptively select the focused patches with a higher $WQ_{SSIM}$ value to form the final fused results. Extensive experiment results have shown that our framework outperforms the state-of-the-art methods in terms of focus information preservation and artifact removal under real-world challenging scenarios.